# A multidimensional measurement of photorealistic avatar quality of experience

### v1

@@ Abstract.
Photorealistic avatars are human avatars that look, move, and talk like real people. The performance of photorealistic avatars has significantly improved recently based on objective metrics such as PSNR, SSIM, LPIPS, FID, and FVD. However, recent photorealistic avatar publications do not provide subjective tests of the avatars to measure human usability factors. We provide an open source test framework to subjectively measure photorealistic avatar performance in ten dimensions: realism, trust, comfortableness using, comfortableness interacting with, appropriateness for work, creepiness, formality, affinity, resemblance to the person, and emotion accuracy. We show that the correlation of nine of these subjective metrics with PSNR, SSIM, LPIPS, FID, and FVD is weak, and moderate for emotion accuracy. The crowdsourced subjective test framework is highly reproducible and accurate when compared to a panel of experts. We analyze a wide range of avatars from photorealistic to cartoon-like and show that some photorealistic avatars are approaching real video performance based on these dimensions. We also find that for avatars above a certain level of realism, eight of these measured dimensions are strongly correlated. In particular, for photorealistic avatars there is a linear relationship between avatar affinity and realism; in other words, there is no uncanny valley effect for photorealistic avatars in the telecommunication scenario. We provide several extensions of this test framework for future work and discuss design implications for telecommunication systems. The test framework is available at https://github.com/microsoft/P.910.
@@ 4.2. Test set
To test the test framework we created a test set consisting of 237 video clips of 19 different avatar models. For Template A the video clips were of people talking. For Template B the video clips were facial emotions: happy, sad, surprised, fear, anger, and disgust. We found that applying Template A to the facial emotion clips resulted in all the clips being rated as creepy, including the real video clips. Therefore we kept the Template A and Template B clips mutually exclusive.
We used 19 different avatars in the test set. The VASA-1 (Xu et al., 2024a), MS1, MS2, MS3, MS4 and MS5 avatars are from Microsoft. External1, External2 and External3 are 3 variations from the same organization. External1 is the AI model, External2 is the computer graphics based avatar and External3 is the set of avatars we created from the iPhone application from the organization. The Unreal [^1] https://www.unrealengine.com/en-US/metahuman avatars are models created using Epic’s Metahuman avatar creator. The Unreal_pro* avatars are animations of Epic’s Metahuman store models created by professionals, while the non-pro Unreal avatars are created by the authors. The EMO (Tian et al., 2024), MegActor (Yang et al., 2024), Make-Your-Anchor (Huang et al., 2024), DisCo (Wang et al., 2024) and Pose2Img (Qian et al., 2021) avatars were scraped from their respective project pages. The Zoom avatar was a screen capture of the avatar in Zoom’s video conferencing application. The Meta avatar is from the Mark Zuckerberg, Lex Fridman interview [^2] https://www.youtube.com/watch?v=MVYrJJNdrEg. Audio2Photoreal and LivePortrait (Guo et al., 2024) were created using their open sourced models on GitHub. 9 of the 19 avatar models have more than one viewpoint variation (frontal, $45^{\circ}$ off-center and $90^{\circ}$ off-center). Thus we have a total of 40 different avatar conditions with an average of 5.4 clips per variation for Template A. This also includes the baseline clips of real recordings of the actors. For Template B, we only used models that had input video clips available, resulting in 26 different avatar conditions from 13 different avatars. Table 5 provides the number of people for each avatar, the number viewpoints, whether speech is included, and if there was a reference video driving the avatar.
Table 5. Details of the avatar test set.
Avatar # people # viewpoints Speech Reference video
Audio2Photoreal 4 5 Y N
DisCo 4 1 N N
EMO 2 1 Y Y
External1 4 3 Y Y
External2 4 3 Y Y
External3 4 3 Y Y
LivePortrait 5 1 Y Y
MegActor 4 1 N N
Make-Your-Anchor 6 1 N N
Meta 2 1 Y N
MS1 4 3 Y Y
MS2 5 3 Y N
MS3 5 4 Y Y
MS4 3 1 Y N
MS5 3 1 Y Y
Pose2Img 4 1 N N
Unreal 6 3 Y Y
Unreal_pro* 8 3 Y N
VASA-1 8 1 Y N
@@ 4.3. Reproducibility
To test the reproducibility from run to run, for each template we repeated our crowdsourcing test five times with a mutually exclusive group of workers, on separate days on Amazon Mechanical Turk. We calculated the Mean Opinion Scores (MOS) per clip and per condition (avatar) and show the correlation between different runs and scales in Tables 7 and 8. The results show a strong correlation between different runs at the clip (average PCC=0.968) and condition levels (average PCC=0.987) and confirm the survey provides reproducible scores.
@@ 5. Results and analysis
Table 7. Pearson correlation between different runs of the reproducibility test in condition level.
@@ 5.2. Correlation analysis for Template A
We observed a strong correlation between the dimensions from Template A (Table 9), on average PCC=0.996. This was surprisingly high, which motivated doing the three versions of the survey in Section 4.1 which all exhibited the same high correlation. However, when we consider only low realism avatars (realism $\leq$ 2) then the dimensions are not all highly correlated as shown in Table 11. But for avatars with realism ¿ 2, the dimensions are very highly correlated with an average PCC=0.999 as shown in Table 10.
@@ 5.4. Correlation with objective metrics
We evaluated the effectiveness of objective metrics in representing subjective scores. In this domain, PSNR (Gonzalez and Woods, 2006), SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) are typically reported in publications. We also analyze FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019), which are less commonly reported. We utilized eight models from the study reported in Section 4.3, where the original video was available (overall 31 clips). We removed the background of the person from the original videos and transformed the videos to match each avatar representation using face landmarks. Finally, we calculated the objective metrics for each video clip and aggregated them for the avatar models. Table 12 reports the Pearson’s correlation coefficient and Kendall’s Tau-b rank correlation coefficient [^3] We report Kendall’s Taub-b as it is more robust to outliers and non-normal distributions compare to Spearman’s rank correlation (Croux and Dehon, 2010) making it more suitable for small sample sizes. (Kendall, 1938) between the objective metrics and the subjective scores from Template A and Template B. The results show a weak correlation between the objective metrics and the subjective scores from Template A, and a moderate correlation with scores from Template B for PSNR, SSIM, and LPIPS. This highlights the necessity of using subjective scores for accessing the quality of experience of photorealistic avatars.
@@ 6. Conclusions and future work
We have provided a test framework to measure the quality of experience of photorealistic avatars, and shown it to be accurate and reproducible (RQ3). We have shown the commonly used objective metrics (PSNR, SSIM, LPIPS, FIV, FVD) are weakly correlated to nine of the subjective metrics we define, and moderately correlated to emotion accuracy (RQ1). For avatars with a realism $>$ 2 MOS we have shown in Table 10 that all of the no-reference metrics (Template A) are highly correlated to each other, which means that for Template A only one metric needs to be measured (e.g., realism), while for Template B both metrics need to be measured; this leads to dimensionality reduction of 10 to 3 for the survey (RQ2). Finally, we have shown in the context of photorealistic avatars in telecommunication scenarios that the uncanny valley effect does not exist, which implies photorealistic avatars should be made as realistic as possible to maximize the quality of experience of the avatar (RQ4).
@@
  • Audio video synchronization is very important for photorealistic avatars and can be measured subjectively and compared with objective metrics such as (Chung and Zisserman, 2017).
@@
  • The avatars should be as realistic as possible, ideally matching the realism from real video. This will maximize trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, while minimizing creepiness.
@@
One limitation of this study is the diversity of the avatars studied. We included 17 avatars which included 13 males, 4 females, 11 Caucasians, 6 Asians, and 1 Black. Ideally, we would have included more avatars with an equal number of males and females and a greater number of and more equality distributed races represented.

### v2

@@ Abstract.
Photorealistic avatars are human avatars that look, move, and talk like real people. The performance of photorealistic avatars has significantly improved recently based on objective metrics such as PSNR, SSIM, LPIPS, FID, and FVD. However, recent photorealistic avatar publications do not provide subjective tests of the avatars to measure human usability factors. We provide an open source test framework to subjectively measure photorealistic avatar performance in ten dimensions: realism, trust, comfortableness using, comfortableness interacting with, appropriateness for work, creepiness, formality, affinity, resemblance to the person, and emotion accuracy. We show that the correlation of nine of these subjective metrics with PSNR, SSIM, LPIPS, FID, and FVD is weak, and moderate for emotion accuracy. The crowdsourced subjective test framework is highly reproducible and accurate when compared to a panel of experts. We analyze a wide range of avatars from photorealistic to cartoon-like and show that some photorealistic avatars are approaching real video performance based on these dimensions. We also find that for avatars above a certain level of realism, eight of these measured dimensions are strongly correlated. This means that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video. In addition, because there is a strong linear relationship between avatar affinity and realism, there is no uncanny valley effect for photorealistic avatars in the telecommunication scenario. We provide several extensions of this test framework for future work and discuss design implications for telecommunication systems. The test framework is available at https://github.com/microsoft/P.910.
@@ 1. Introduction
  • RQ5: What avatar performance is needed to have the same level of performance as real video?
@@
  • We show that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video.
@@ 4.2. Test set
To test the test framework we created a test set consisting of 237 video clips of 19 different avatar models. For Template A the video clips were of people talking. For Template B the video clips were also include facial emotions: happy, sad, surprised, fear, anger, and disgust. We found that applying Template A to the facial emotion clips resulted in all the clips being rated as creepy, including the real video clips.
We used 19 different avatars in the test set. The VASA-1 (Xu et al., 2024a), MS1, MS2, MS3, MS4 and MS5 avatars are from Microsoft. External1, External2, and External3 are 3 variations from the same organization. External1 is the AI model, External2 is the computer graphics based avatar and External3 is the set of avatars we created from the iPhone application from the organization. The Unreal [^1] https://www.unrealengine.com/en-US/metahuman avatars are models created using Epic’s Metahuman avatar creator. The Unreal_pro avatars are animations of Epic’s Metahuman store models created by professionals, while the non-pro Unreal avatars are created by the authors. The EMO (Tian et al., 2024), MegActor (Yang et al., 2024), Make-Your-Anchor (Huang et al., 2024), DisCo (Wang et al., 2024), Pose2Img (Qian et al., 2021), and Meta2 (Cao et al., 2022) avatars were scraped from their respective project pages. The Zoom avatar was a screen capture of the avatar in Zoom’s video conferencing application. The Meta1 avatar is from the Mark Zuckerberg, Lex Fridman interview [^2] https://www.youtube.com/watch?v=MVYrJJNdrEg. Audio2Photoreal and LivePortrait (Guo et al., 2024) were created using their open sourced models on GitHub. 9 of the 19 avatar models have more than one viewpoint variation (frontal, $45^{\circ}$ off-center and $90^{\circ}$ off-center). Thus we have a total of 40 different avatar conditions with an average of 5.4 clips per variation for Template A. This also includes the baseline clips of real recordings of the actors. For Template B, we only used models that had input video clips available, resulting in 26 different avatar conditions from 13 different avatars with an average of 5 clips per variation. Table 5 provides the number of people for each avatar, the number viewpoints, whether speech is included, and if there was a reference video driving the avatar.
Table 5. Details of the avatar test set. Races are Caucasian, Asian, Black, and Genders are Male/Female.
Avatar # people Race C/A/B Gender M/F # viewpoints Speech Reference video
Audio2Photoreal 4 3/0/1 2/2 5 Y N
DisCo 4 2/2/0 3/1 1 N N
EMO 2 1/1/0 1/1 1 Y Y
External1 4 3/0/1 3/1 3 Y Y
External2 4 3/0/1 3/1 3 Y Y
External3 4 2/2/0 4/0 3 Y Y
LivePortrait 4 2/2/0 4/0 1 Y Y
MegActor 4 2/1/1 2/2 1 N N
Make-Your-Anchor 6 3/3/0 5/1 1 N N
Meta1 2 2/0/0 2/0 1 Y N
Meta2 6 3/2/1 4/2 1 N Y
MS1 4 2/2/0 4/0 3 Y Y
MS2 5 3/2/0 3/2 3 Y N
MS3 5 4/1/0 4/1 4 Y Y
MS4 4 1/3/0 4/0 1 Y N
MS5 2 2/0/0 2/0 1 Y Y
Pose2Img 4 2/2/0 3/1 1 N N
Unreal 4 2/2/0 4/0 3 Y Y
Unreal_pro 10 5/1/4 7/3 3 Y N
VASA-1 6 4/1/1 3/3 1 Y N
Total 88 51/27/10 67/21
@@ 4.3. Reproducibility
To test the reproducibility from run to run, for each template we repeated our crowdsourcing test five times with a mutually exclusive group of workers, on separate days on Amazon Mechanical Turk. On average we have collected 22.8 votes per video clip using Template A and 22.4 votes using Template B in each run. We calculated the Mean Opinion Scores (MOS) per clip and per condition (avatar variation) and show the correlation between different runs and scales in Tables 7 and 8. The results show a strong correlation between different runs at the clip (average PCC=0.968) and condition levels (average PCC=0.987) and confirm the survey provides reproducible scores.
@@ 5. Results and analysis
Table 7. Pearson correlation between different runs of the reproducibility test in model level.
@@ 5.2. Correlation analysis for Template A
We observed a strong correlation between the dimensions from Template A (Table 9), on average PCC=0.996. This was surprisingly high, which motivated doing the three versions of the survey in Section 4.1 which all exhibited the same high correlation. However, when we consider only low realism avatars (realism $\leq$ 2) then the dimensions are not all highly correlated as shown in Table 11. But for avatars with realism ¿ 2, the dimensions are very highly correlated with an average PCC=0.999 as shown in Table 10. One important implication of this high correlation is that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video.
@@ 5.4. Correlation with objective metrics
We evaluated the effectiveness of objective metrics in representing subjective scores. In this domain, PSNR (Gonzalez and Woods, 2006), SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) are typically reported in publications. We also analyze FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019), which are less commonly reported. We utilized eight models from the study reported in Section 4.3, where the original video was available (overall 31 clips). We removed the background of the person from the original videos and transformed the videos to match each avatar representation using face landmarks. Finally, we calculated the objective metrics for each video clip and aggregated them for the avatar models. Table 12 reports the Pearson’s correlation coefficient and Kendall’s Tau-b rank correlation coefficient [^3] We report Kendall’s Tau-b as it is more robust to outliers and non-normal distributions compare to Spearman’s rank correlation (Croux and Dehon, 2010) making it more suitable for small sample sizes. (Kendall, 1938) between the objective metrics and the subjective scores from Template A and Template B. The results show a weak correlation between the objective metrics and the subjective scores from Template A, and a moderate correlation with scores from Template B for PSNR, SSIM, and LPIPS. This highlights the necessity of using subjective scores for accessing the quality of experience of photorealistic avatars.
@@ 6. Conclusions and future work
We have provided a test framework to measure the quality of experience of photorealistic avatars, and shown it to be accurate and reproducible (RQ3). We have shown the commonly used objective metrics (PSNR, SSIM, LPIPS, FIV, FVD) are weakly correlated to nine of the subjective metrics we define, and moderately correlated to emotion accuracy (RQ1). For avatars with a realism $>$ 2 MOS we have shown in Table 10 that all of the no-reference metrics (Template A) are highly correlated to each other, which means that for Template A only one metric needs to be measured (e.g., realism), while for Template B both metrics need to be measured; this leads to dimensionality reduction of 10 to 3 for the survey (RQ2). We have shown in the context of photorealistic avatars in telecommunication scenarios that the uncanny valley effect does not exist, which implies photorealistic avatars should be made as realistic as possible to maximize the quality of experience of the avatar (RQ4). Finally, we show that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video (RQ5).
@@
  • To better understand what areas to improve for realism, we can ask the rater to fill in why they gave it that rating and provide binary options such as (1) The avatar was distorted, (2) The lip sync was off, (3) Absence of microdetails and imperfections, (4) Inaccurate lighting and shadows, (5) Unnatural textures and materials, and (6) Other (the rater provides a description).
@@
  • The avatars should be as realistic as possible, ideally matching the realism from real video. This will maximize trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, while minimizing creepiness. Critically, avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video.
@@
One limitation of this study is the diversity of the avatars studied. We included 20 avatars which included 67 males, 21 females, 51 Caucasians, 27 Asians, and 10 Blacks. Ideally, we would have included more avatars with an equal number of males and females and a greater number of and more equality distributed races represented.